### v2

# On the PM2.5 – Mortality Relationship: A Bayesian Model for Spatio-Temporal Confounding

Abstract
In epidemiological studies about the effects of air pollution on public health, it is common practice to incorporate an arbitrary smooth function of time (or space) in time-series (or spatial) data to account for unknown confounding effects. These smooth functions are inherently linked to exposure, thereby increasing the difficulty of assessing the exposure effect (i.e. the concurvity problem). This issue is particularly challenging in spatio-temporal analyses, where it has been less explored so far. To study the effects of fine particulate matter on mortality in Italy, a model that simultaneously accounts for spatio-temporal confounding and for the non-linear form of the effect of interest is needed. Henceforth, we propose a Bayesian spatial dynamic generalized linear model, which allows for non-linear associations and for a decomposition of the exposure effect. This decomposition accommodates associations with the outcome at fine and coarse temporal and spatial scales of variation. These features, when combined, allow reducing the spatio-temporal confounding bias and recovering the shape of the association, as demonstrated through simulation studies. The real-data analysis indicates that the exposure effect shows a clear temporal dynamic, with peaks during summer. We argue that this may be due to interactions of particulate matter concentrations with air temperature and unmeasured confounders.
Keywords: Air pollution, Bayesian, Concurvity, Dynamic generalized linear model, Non-linearity, Spatio-temporal confounding.
1 Introduction
Air pollution is a major environmental risk factor and undermines public health, causing a wide range of health problems, including premature deaths. To date, there is a very rich literature demonstrating the short- and long-term consequences of air pollution on health outcomes (see, e.g., Dominici et al. (2002), Janes et al. (2007), Zeger et al. (2008), Greven et al. (2011), Dai et al. (2014), Di et al. (2017), Wu et al. (2020), Brunekreef et al. (2021b), Renzi et al. (2021), Dominici et al. (2022), Gamerman et al. (2022), Requia et al. (2024)).
Within the ecological regression framework, one of the challenges encountered is the determination of the shape of the exposure–outcome relationship, which is commonly assumed to be constant and linear. We say that the relationship is linear if the exposure enters linearly in the specification of the link function. However, if this assumption is violated, misspecification bias would result. We acknowledge that all models are inherently flawed, but some of them are so misspecified that they can lead to incorrect inferential statements (Gelman and Shalizi, 2013).
In this paper, the ultimate goal is to correctly evaluate the short-term association between particulate matter with a diameter less than $2.5\mu m$ (PM2.5) and all-cause mortality. The data span the period between 2018 and 2022 and are available for each health district of two Italian regions (see Section 2). We argue that the commonly adopted assumption of a constant and linear effect, without considering interaction with confounding variables, may be inadequate, given the spatial and temporal variability of PM2.5 and its associated health risks. Specifically: (i) the physical properties and chemical composition of the PM mixture exhibits substantial variation across locations and over time, driven by environmental conditions, human activities, and episodic phenomena such as sandstorms and wildfires (Chen et al., 2013, Kelly and Fussell, 2012, Masselot et al., 2022, Peng et al., 2005, Zhou et al., 2021); (ii) human susceptibility to PM2.5 is not static, but rather modulated by temporally and spatially varying factors, including extreme temperatures, which can exacerbate health risks in specific regions or seasons (Bergmann et al., 2020, Peng et al., 2005, Roberts, 2004, Li et al., 2017); (iii) the relationship between true and measured exposure varies across time and space, with seasonal factors like increased summer ventilation potentially reducing exposure measurement error (Bergmann et al., 2020, Peng et al., 2006, Roberts, 2004, Stafoggia et al., 2008).
To account for the potential non-linearity of the exposure–outcome relationship, various strategies are available. An exposure-response function is often estimated (see, among others, Brunekreef et al., 2021b, Cork et al., 2023, Dominici et al., 2022, Renzi et al., 2021). Alternatively, it would be appropriate to specify a model that incorporates the inherent spatial and temporal variations of the exposure effect (Gelfand et al., 2003). By doing so, we can analyze whether the strength of the relationship between PM2.5 and mortality have changed over different time periods and across health districts. The temporal dynamic of the exposure effect allows investigating the influence of various factors that have emerged since 2020, such as lockdowns, the use of facial masks, physical distancing and social restrictions. Indeed, these measures have significantly altered people’s habits and interactions with the environment. Furthermore, the health districts are characterized, for example, by morphological and climatic differences which may influence the exposure–outcome relationship. Therefore, it is important to allow for spatially varying coefficients in the model.
Time-varying coefficients models (TVCMs) have been previously considered in the literature. For example, the exposure effect is modeled either as a periodic function (Chen et al., 2013, Peng et al., 2005) or as a season-specific step function (Dai et al., 2014, Klompmaker et al., 2023). Moreover, Chiogna and Gaetan (2002) and Lee and Shaddick (2008) employ dynamic generalized linear models (DGLMs; West and Harrison, 1997) when analyzing the relationship between air pollution and mortality with time-series data.
Spatio-temporally varying coefficients models (STVCMs) are reviewed by Banerjee et al. (2014) and Sahu (2022), among others, and have been applied to various fields, including healthcare (Cai et al., 2013, Song et al., 2020), air pollution and climate studies (Paez et al., 2008, Bakar et al., 2015). This class of models provide a flexible framework to capture smooth space-time variations of the exposure–outcome association.
A second major challenge in ecological regression models relates to the problem of confounding. When data are spatially correlated, this problem has received a lot of attention in recent years — see, e.g., Reich et al. (2006), Paciorek (2010), Thaden and Kneib (2018), Dupont et al. (2022), Marques et al. (2022), Azevedo et al. (2023) and Zaccardi et al. (2025). Gilbert et al. (2023) and Khan and Berrett (2023) argue that multiple definitions of “spatial confounding” co-exist in the literature. In this paper, we extend one of their notions and refer to spatio-temporal confounding as the problematic assessment of the exposure effect due to concurvity, that is the correlation between exposure and an unknown smooth function of space-time that affects the outcome (Gilbert et al., 2023, Peng et al., 2006). In other words, factors that vary on similar spatial and temporal scales as the exposure are considered potential confounders. Although information on some of these confounders is available, such as meteorological variables, it is likely that others are unmeasured or difficult to quantify, such as influenza epidemics or changes in population structure (Peng et al., 2006). Not correctly accounting for unmeasured confounding may lead to an incorrect evaluation of the exposure–outcome relationship (i.e., confounding bias), as the model may not be able to distinguish between the effect of the exposure and that of the confounders.
Despite the complexity of this problem, just few authors have recently proposed approaches that explicitly account for spatio-temporal confounding. For instance, Reich et al. (2021) discuss extensions of causal inference methods. Other works have adopted a “data analysis” perspective (Khan and Berrett, 2023), where confounding results from the collinearity between exposure and structured random effects. Within this strand of research, Prates et al. (2022) propose the transformed multivariate Gaussian Markov random field model (TMGMRF), an extension of the approach by Prates et al. (2015), where Poisson intensities or Bernoulli rates are modeled directly and without the need of a link function. A copula structure captures the spatio-temporal dependence and guarantees the disentanglement between fixed and random effects, hence their model should be able to control for confounding bias. A spatio-temporal linear model is considered by Adin et al. (2023), where the random effects are orthogonalized with respect to the fixed effects: in other words, this work extends the restricted spatial regression model (RSR; Hodges and Reich, 2010, Reich et al., 2006). These approaches, however, are not intended to adjust for unmeasured confounders (Khan and Calder, 2022).
A comprehensive data-driven approach that can capture the temporal dynamics of the exposure effect while simultaneously controlling for spatio-temporal confounding is still lacking. To tackle both problems, we propose, under the Bayesian paradigm, the spatial DGLM for confounding (SDGLMC), which has several advantages. First, spacetime-varying intercept and slope coefficient are introduced in the model specification to account for unmeasured spatio-temporal confounders and non-linear exposure–outcome associations, respectively. Second, since DGLMs are only locally linear, they can approximate any smooth function of time (Gamerman and Lopes, 2006, Granger, 2008), hence the proposed model should not be affected by misspecification bias. Third, SDGLMC can be seen as a generalization of other methods in the literature. For example, the approaches by Chen et al. (2013) and Peng et al. (2005) cannot account for exceptional events such as the outbreak of the SARS-CoV-2 pandemic. The models proposed, among others, by Dai et al. (2014) and Klompmaker et al. (2023) assume that the effect varies between seasons, but is constant within each season. In other terms, the latter works implicitly assume the possibility of abrupt changes in the exposure effect between adjacent seasons, however these are unlikely in practice, because the exposure effect depends, e.g., on meteorological conditions and on human activity patterns that change gradually over time (Peng et al., 2005, Bhaskaran et al., 2013). In contrast, the proposed model can overcome both drawbacks, thanks to the ability of DGLMs to capture gradual changes and to emphasize both seasonal and non-systematic variations (Lütkepohl, 2005, Petris et al., 2009, Bitto and Frühwirth-Schnatter, 2019). Additionally, SDGLMC extends the approaches by Paez et al. (2008), Cai et al. (2013) and Bakar et al. (2015) since it includes a space-time interaction term in the intercept, thus enabling control for complex spatio-temporal confounding patterns.
As a fourth major advantage, following Greven et al. (2011) and Janes et al. (2007), the exposure effect is decomposed into two orthogonal components, each describing the (non-linear) association with the health outcome at a different scale of variation (see Section 3). The focus of inference is the component that captures the association at a finer scale — where confounding by unmeasured factors is less likely (Greven et al., 2011) — which is assumed to exhibit spatio-temporal dynamics. Furthermore, SDGLMC provides explicit interpretation of each part of these dynamics through an analysis-of-variance-like decomposition (see Section 3.1).
This paper is organized as follows. Section 2 offers a brief description of the data that motivated this study. Sections 3 and 4 provide detailed information on the proposed model and on the inferential procedure, respectively. Section 5 presents existing approaches from the literature to which the results of the proposed model are compared. In Section 6, a simulation study is conducted to investigate the ability of SDGLMC in controlling for spatio-temporal confounding when the true association is known. The results from the real-data application are analyzed in Section 7. Finally, Section 8 provides a discussion.
2 The Data
This section is dedicated to the description of the dataset that motivated our research, specifically we analyze the association between daily PM2.5 concentrations and all-cause mortality counts among the elderly (i.e. people aged 65 or more) in Piemonte and Lombardia regions (Italy) between 2018 and 2022.
The data are retrieved from publicly available sources, and full details are given in Section A of the Supplementary Material. Mortality counts are collected from the Italian National Institute of Statistics (ISTAT, 2023a), and are observed on an irregular lattice comprising ${N}=117$ non-overlapping areal units termed health districts (HDs). The study period comprises ${T}=1826$ consecutive days, hence $Y_{it}$ denotes the number of deaths (by any cause) occurred in district $i=1,\dots,{N}$ during day $t=1,\dots,{T}$. The PM2.5 concentrations are used as a measure of the exposure in district $i$ during day $t$, denoted as $X_{it}$, and are collected from the Copernicus Atmosphere Monitoring Service’s (CAMS) Atmosphere Data Store (INERIS et al., 2022, CAMS, 2023). We also acquire information on temperature and relative humidity, which are potential confounders, from the ERA5 reanalyses dataset available at the Copernicus Climate Change Service’s (C3S) Climate Data Store (C3S, 2023, Hersbach et al., 2018). The remote-sensed measurements are spatially and temporally aggregated to match the resolution of the health data.
A measure usually employed for exploratory analyses is the standardized mortality ratio (SMR), which is calculated as $SMR_{it}=Y_{it}/E_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$. The denominator, $E_{it}$, denotes the expected number of deaths, and is computed allowing for the underlying age–sex stratification of the population within each HD (see also Section A of the Supplementary Material). Figure 1 (left panel) depicts the average daily SMR for each HD over the years 2018–2022, whose mean value is $1.25$ over the whole study domain, and spatial clustering is noticeable. The right panel shows the long-term average PM2.5 concentrations for each HD, and on average the Lombardia region is more polluted than Piemonte ($19.41\,\mu g/m^{3}$ and $15.18\,\mu g/m^{3}$, respectively). Other descriptive statistics are reported in Section B of the Supplementary Material.
Figure 1: Long-term averages over the whole study period (2018–2022) of the SMR due to all-cause mortality (left) and of the PM2.5 levels (right) for each HD. The thicker solid line separates Piemonte (the westernmost region) from Lombardia (the easternmost region).
Figure 2 shows the average daily SMR (on the logarithmic scale) for Lombardia and Piemonte. The seasonal pattern is similar in both regions, with the lowest values in summer, except for the year 2020. Two peaks can be noticed, where the log(SMR) more than doubled: the first one occurred in April 2020 and the other one in November 2020. These are due to the first and second “waves” of the SARS-CoV-2 pandemic, respectively. The first wave was more severe in Lombardia, while the second one was more severe in Piemonte. In addition, the two peaks in the SMR in 2020 are not reflected in the PM2.5 time series. As can be noticed from Table B.3 in the Supplementary Material, the overall 2020 annual mean of PM2.5 levels was the lowest among the five years considered in our analysis.
Figure 2: Time series of the average daily SMR (on the log scale, orange lines) and average daily PM2.5 concentrations (black lines) for Piemonte (top) and Lombardia (bottom) regions. The vertical dashed lines separate each calendar year. Blue vertical bands correspond to the national lockdown periods.
Figure 2 also highlights other two aspects. One is that the highest values of the SMR in 2018 were reached at the beginning of January, suggesting that the influenza epidemic reached its peak before the end of 2017. This is different from the year 2019, where the SMR peaks were reached in February. The other aspect is that there was an increase in mortality during summer 2022. Possible reasons may be the occurrence of prolonged heat waves (indeed, the average daily temperature time-series has its absolute maximum in July 2022), or the outbreak of an influenza epidemic.
However, to study the association of interest, it is necessary to account for the effects of measured and unmeasured confounders. Failing to do so may lead to incorrect inferences, due to concurvity. The proposed approach, which is detailed in the next section, is framed to overcome this issue, among others.
3 The Proposed Model
It is assumed that the outcome, $Y_{it}$, follows a Poisson distribution with mean $\mu_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$. The linear predictor, $\vartheta_{it}=\ln(\mu_{it})$, is modeled as a function of exposure, $X_{it}$, measured confounders, $\mathbf{M}_{it}=({M}_{1it},\dots,{M}_{{p_{m}}it})^{\prime}$, and an unmeasured confounder, $Z_{it}$, which is a smooth function of space and time. Assuming that ${\mathbf{M}}_{it}$ and $Z_{it}$ are spatially and temporally smoother than the exposure, the model can be specified as follows:
$Y_{it}|\mu_{it}$ $\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{P}oi(\mu_{it})$ (1)
$\vartheta_{it}$ $={\color[rgb]{0,0,0}\ln(\mu_{it})}=o_{it}+\beta_{0}+f(X_{it};\bm{\gamma}_{f})+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+Z_{it}+u_{it}\,,$ (2)
where $o_{it}=\log(E_{it})$ is the offset, $\beta_{0}$ is an intercept, $f(\cdot;\cdot)$ is an unknown smooth function parameterized by the vector $\bm{\gamma}_{f}$, $\bm{\alpha}$ is a ${p_{m}}$-dimensional vector of coefficients, and $u_{it}$ is a spatially and temporally uncorrelated Gaussian random variable, assumed to have zero mean and constant variance, i.e. $u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})$.
If all confounders were measured and $f(\cdot;\cdot)$ was modeled using basis functions, the model specification in Equations (1)–(2) would belong to the family of generalized additive models (GAMs; see Wood, 2017). Alternatively, considering that $f(\cdot;\cdot)$ can be adequately approximated by using the first-order Taylor’s expansion, we have:
$f(X_{it};\bm{\gamma}_{f})=f(\widehat{X}_{it};\bm{\gamma}_{f})+f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})+\varpi_{it}\,,$ (3)
where $\widehat{X}_{it}$ denotes the large-scale (or overall) spatio-temporal trend, $f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})$ is the first derivative of $f(\cdot;\bm{\gamma}_{f})$ with respect to its first argument, evaluated at $\widehat{X}_{it}$, and $\varpi_{it}$ denotes the Taylor’s remainder. To calculate the spatio-temporal trend, $X_{it}$ can be regressed on a smooth function of space and time (see also Janes et al., 2007):
${\color[rgb]{0,0,0}X_{it}=h({\bf s}_{i},t;df)+\nu_{it}\,,}$ (4)
where ${\bf s}_{i}$ denotes the spatial coordinates of the centroid of areal unit $i$, $df$ are the degrees of freedom of the smooth term, and $\nu_{it}$ is an uncorrelated Gaussian random variable with zero mean and constant variance. The overall trend is then defined as the fitted values obtained from the model (4), i.e. $\widehat{X}_{it}=\widehat{h}({\bf s}_{i},t;df)$.
Equation (3) provides a decomposition of the exposure effect into two components describing large and small (or local) scales of variation. These components are functions of $\widehat{X}_{it}$ and $X_{it}-\widehat{X}_{it}$, respectively. They are orthogonal because $X_{it}-\widehat{X}_{it}$ is the residual term from the regression in Equation (4). Henceforth, the decomposition above should help to disentangle the effect of interest from that of unmeasured confounding factors. The rationale for using detrended exposure, as discussed in the econometric and spatio-temporal literature, is to remove long-term and large-scale spatial variation, preserving only short-term and local fluctuations. This helps reduce the risk of spurious associations driven by underlying trends (Wooldridge, 2020). From an epidemiological point of view, consider that $\widehat{X}_{it}$ captures trend and seasonal components of the exposure, therefore it is expected to be highly correlated with $Z_{it}$, given that the latter is smooth, while the correlation between $X_{it}-\widehat{X}_{it}$ and $Z_{it}$ should be close to zero. Therefore, the unmeasured confounder is more likely to affect the association between outcome and exposure at large scales (Greven et al., 2011, Janes et al., 2007), so the first term in the decomposition of $f(X_{it};\bm{\gamma}_{f})$ in Equation (3) might be biased, but it is less likely that the second term, $f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})$, is affected by confounding bias.
By decomposing $f(X_{it};\bm{\gamma}_{f})$ in Equation (2), we thus have:
$\vartheta_{it}=o_{it}+\beta_{0}+f(\widehat{X}_{it};\bm{\gamma}_{f})+f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+Z_{it}+\tilde{u}_{it}\,.$ (5)
where $\tilde{u}_{it}=u_{it}+\varpi_{it}$ is such that $\tilde{u}_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tilde{\tau}_{u}^{2})$.
For each $i$, the unknown function evaluated at the overall trend of the exposure, $f(\widehat{X}_{it};\bm{\gamma}_{f})$, and its first derivative in Equation (5) can be seen as functions of time, hence time-varying coefficients can adequately represent them (see also West and Harrison, 1997, Section 3.1). Therefore, we propose to replace the linear predictor in Equation (2) with:
$\vartheta_{it}=o_{it}+\beta_{0it}+\beta_{1it}(X_{it}-\widehat{X}_{it})+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+\tilde{u}_{it}\,,$ (6)
where $\beta_{1it}=f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})$ is the spacetime-varying coefficient representing the local association between exposure and outcome (in the sense of Janes et al. (2007)), while $\beta_{0it}$ is a spacetime-varying intercept,
$\beta_{0it}=\beta_{0}+f(\widehat{X}_{it};\bm{\gamma}_{f})+Z_{it}\,,$
that should account for both the large-scale exposure effect and the unmeasured confounder.
We now provide two examples of the function $f(\cdot;\cdot)$ and of the resulting varying coefficients in Equation (6). These show that our model generalizes other methods in the literature.
Example 3.1.
If $f(\cdot;\bm{\gamma}_{f})$ is a constant function, i.e. $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}$, then $\beta_{1it}=0$ and $\beta_{0it}=\beta_{0}+\gamma_{f0}+Z_{it}$.
Example 3.2.
If $f(\cdot;\bm{\gamma}_{f})$ is a linear function, i.e. $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}$, then $\beta_{1it}=\gamma_{f1}$ and $\beta_{0it}=\beta_{0}+\gamma_{f0}+\gamma_{f1}\widehat{X}_{it}+Z_{it}$. The model proposed by Greven et al. (2011) can be cast in this case, with $\gamma_{f0}=0$. The authors, however, assume further that the intercepts do not depend on time.
3.1 Specification of the Spacetime-Varying Coefficients
Although it is possible to model $\beta_{0it}$ and $\beta_{1it}$ directly, we decompose these spacetime-varying coefficients as follows to enhance prior elicitation and interpretability (see also Bakar et al., 2015, Cai et al., 2013, Paez et al., 2008):
$\beta_{kit}=\overline{\delta}_{k}+\delta_{ki}+\delta_{kt}+\delta_{kit}^{*}\,,\quad k=0,1\,,$
where $\overline{\delta}_{k}$ is the global effect of the $k$th covariate over the whole study period, $\delta_{ki}$ is the area-specific effect of the $k$th covariate, $\delta_{kt}$ is the time-specific effect of the $k$th covariate, and $\delta_{kit}^{*}$ is the spatio-temporal interaction effect of the $k$th covariate. For each $i$, $\delta_{ki}$ can be interpreted as the average (across time) spatial deviation from the global effect. In particular, $\delta_{1i}$ captures time-independent spatial heterogeneity in the exposure–outcome association at the local scale of variation (recall that the regressor is $X_{it}-\widehat{X}_{it}$). It can inform about where populations are most vulnerable to the effects of air pollution.
For each time instant, $\delta_{kt}$ can be interpreted as the average (across space) temporal changes from the global effect. Hence, $\delta_{1t}$ represents the temporal evolution of the association of interest, and highlights when peaks can occur. The interaction effect, $\delta_{kit}^{*}$, captures the area- and time-specific deviations from the global effect that are left after accounting for the temporal and spatial effects. The simulation scenarios in Section 6 were built based on possible values of $\beta_{1it}$ that are very small in real applications. Hence, given the tiny contribution of the term $\beta_{1it}(X_{it}-\widehat{X}_{it})$ to the augmented-data likelihood (see Section 4.2), it is difficult to recover the interaction effect of the exposure, even in the absence of confounding. Therefore, we assume that $\delta_{1it}^{*}$ is zero. This assumption is in line with previous studies (see, e.g., Bakar et al., 2015, Paez et al., 2008).
In SDGLMC, we implement a DGLM with random walk (RW) processes to model the temporal effects, $\delta_{0t}$ and $\delta_{1t}$, and the interaction effect, $\delta_{0it}^{*}$. We propose the following state-space formulation for the latent process and the regression coefficients, for $t=1,\dots,{T}$:
$\bm{\vartheta}_{t}$ $=\mathbf{o}_{t}+{\bf 1}_{N}(\overline{\delta}_{0}+\delta_{0t})+\tilde{\bm{\delta}}_{0}+\bm{\delta}_{0t}^{*}+\mathbf{\tilde{X}}_{t}(\overline{\delta}_{1}+\delta_{1t})+\mathbf{\tilde{X}}_{t}\odot\tilde{\bm{\delta}}_{1}+{\mathbf{M}}_{t}\bm{\alpha}+\tilde{{\bf u}}_{t}$ (7)
$\delta_{0t}$ $=\delta_{0,t-1}+w_{0t},\quad w_{0t}\sim\mathcal{N}(0,\sigma_{w0}^{2})$ (8)
$\bm{\delta}_{0t}^{*}$ $=\bm{\delta}_{0,t-1}^{*}+\mathbf{w}_{0t}^{*},\quad\mathbf{w}_{0t}^{*}\sim ICAR(\sigma_{w0}^{*2})$ (9)
$\delta_{1t}$ $=\delta_{1,t-1}+w_{1t},\quad w_{1t}\sim\mathcal{N}(0,\sigma_{w1}^{2})$ (10)
where the symbol $\odot$ denotes the Hadamard product, $\bm{\vartheta}_{t}=(\vartheta_{1t},\dots,\vartheta_{{N}t})^{\prime}$, $\mathbf{o}_{t}=(o_{1t},\dots,o_{{N}t})^{\prime}$, $\tilde{\bm{\delta}}_{k}=(\delta_{k1},\dots,\delta_{k{N}})^{\prime}$ are the spatial-specific effects for $k=0,1$, $\tilde{{\bf u}}_{t}=(\tilde{u}_{1t},\dots,\tilde{u}_{{N}t})^{\prime}\sim\mathcal{N}({\bf 0},\tilde{\tau}_{u}^{2}{\bf I}_{N})$, $\bm{\delta}_{0t}^{*}=(\delta_{01t}^{*},\dots,\delta_{0{N}t}^{*})^{\prime}$, and the observation matrices are as follows:
$\mathbf{\tilde{X}}_{t}=\begin{bmatrix}X_{1t}-\widehat{X}_{1t}\\ \vdots\\ X_{{N}t}-\widehat{X}_{{N}t}\end{bmatrix}\quad\text{and}\quad\mathbf{M}_{t}=\begin{bmatrix}\mathbf{M}_{1t}^{\prime}\\ \vdots\\ \mathbf{M}_{{N}t}^{\prime}\end{bmatrix}\,.$
Note that (9) is one of five alternatives that will be discussed further in Section 7. Similarly to $\mathbf{w}_{0t}^{*}$, to capture the spatial correlation among the area-specific effects, $\tilde{\bm{\delta}}_{0}$ and $\tilde{\bm{\delta}}_{1}$ are modeled as intrinsic conditional autoregressive (ICAR; see, e.g., Banerjee et al., 2014) processes, denoted as $ICAR(\sigma_{\delta k}^{2})$, with density proportional to
$p(\tilde{\bm{\delta}}_{k})\propto\exp\biggl{\{}-\frac{1}{2\sigma_{\delta k}^{2}}\tilde{\bm{\delta}}_{k}^{\prime}(\mathbf{D}-\mathbf{W})\tilde{\bm{\delta}}_{k}\biggr{\}},\quad k=0,1\,,$
where $\mathbf{W}$ is the spatial adjacency matrix with elements $(\mathbf{W})_{ij}=1$ if areal units $i$ and $j$ are neighbors, and 0 otherwise, $\mathbf{D}$ is a diagonal matrix with the sum of the rows of $\mathbf{W}$, and $\sigma_{\delta k}^{2}$ is a variance parameter. Such structure has been previously considered by Franco-Villoria et al. (2019, 2022) to model spatially varying coefficients.
4 Bayesian Inference and Computations in SDGLMC
The first stages of the Bayesian hierarchical model are defined by Equations (1) and (7)–(10). To complete it, a prior distribution must be assigned on the parameter set, $\bm{\Theta}=\{\overline{\delta}_{0},\delta_{00},\bm{\delta}_{00}^{*},\overline{\delta}_{1},\delta_{10},\bm{\alpha},\sigma_{w0}^{2},\sigma_{\delta 0}^{2},\sigma_{w0}^{*2},\sigma_{w1}^{2},\sigma_{\delta 1}^{2},\tilde{\tau}_{u}^{2}\}$.
4.1 Prior Specification
Independent prior distributions are imposed on each parameter in $\bm{\Theta}$. As for the time-varying coefficients, distributional assumptions are required on the initial states (i.e., when $t=0$), and on the variances of innovation errors. We impose that $\delta_{k0}\sim\mathcal{N}(0,V_{\delta k})$, $k=0,1$, where $V_{\delta 0}=4$ and $V_{\delta 1}=4\times 10^{-6}$. Also, $\bm{\delta}_{00}^{*}\sim\mathcal{N}(\mathbf{0},\mathbf{V}_{\delta}^{*})$, with $\mathbf{V}_{\delta}^{*}=10\mathbf{I}_{N}$.
Inverse-gamma priors are imposed on $\sigma_{w0}^{*2}\sim\mathcal{IG}(a_{w0}^{*},b_{w0}^{*})$, and similarly on $\sigma_{wk}^{2}\sim\mathcal{IG}(a_{wk},b_{wk})$ and $\sigma_{\delta k}^{2}\sim\mathcal{IG}(a_{\delta k},b_{\delta k})$, for $k=0,1$. For $k=0$, all the hyperparameters are assumed to be $0.01$. For $k=1$, choosing the hyperparameters requires more care. Recall that capturing gradual changes is a desired feature of the proposed model, and the variances $\sigma_{wk}^{2}$ and $\sigma_{\delta 1}^{2}$ play a crucial role to this end. For instance, if their prior means are “large”, the variations in the temporal and spatial effects, respectively, become unlikely pronounced, whereas if their prior means are “small”, the “true” variation is captured only in part. The definition of “large” and “small” mainly depends on the sample size and on how much the prior dominates the likelihood.
Therefore, several values for the hyperparameters in both inverse-gamma distributions are considered to evaluate the sensitivity of the posterior distribution to the prior. In the case of both simulated and real data (see Sections 6 and 7, respectively), we find that $a_{w1}=10$, $b_{w1}=10^{-3}$, and $a_{\delta 1}=1,b_{\delta 1}=0.01$ provide a good balance between the prior and the likelihood.
Finally, the global effects, $\overline{\delta}_{0}$ and $\overline{\delta}_{1}$, and the coefficients of the measured confounders, $\bm{\alpha}$, are assumed independently and normally distributed with mean zero and variance $10^{6}$. Following Wakefield (2007), it is assumed that the residual relative risks, $\tilde{u}_{it}$, follow a log-t distribution with two degrees of freedom, and with $95\%$ of prior mass within the interval $(0.98,1.02)$, hence we obtain the prior $\tilde{\tau}_{u}^{2}\sim\mathcal{IG}(1,2.1\times 10^{-5})$.
4.2 Computations
As a first step of the analysis, the overall temporal trend, $\widehat{X}_{it}$, is obtained as the fitted values from the model in Equation (4), where $h({\bf s}_{i},t;df)$ represents a principal spline of space and time (see, e.g., Mardia et al., 1998, Sahu and Mardia, 2005, Fontanella et al., 2013) with $df$ bases. This two-step strategy should be preferred to a joint modeling of the exposure and the outcome, as it allows avoiding feedback from the outcome to the exposure that may lead to biased estimators of the exposure effect (see, e.g., Stephens et al., 2023).
An MCMC scheme is then implemented to simulate from the posterior distribution of $\bm{\Theta}$. The algorithm proposed by Chan and Jeliazkov (2009) is used to sample from the full conditionals of the time-varying coefficients. It is an efficient simulation algorithm that allows to sample all the regression coefficients (for all the time instants) in a single step. The remaining full conditional distributions are inverse-gamma distributions. The latent process $\vartheta_{it}$ is an exception and is sampled using Metropolis-Hastings steps. To avoid identifiability issues, all time-specific, area-specific and interaction effects are centered at zero during each MCMC iteration (see, e.g., Knorr-Held, 2000, Franco-Villoria et al., 2022). Further details can be found in Section C of the Supplementary Material.
5 Competing Approaches
In this section, we briefly describe the competing approaches that are considered in the simulation study (in Section 6) and in the real data application (in Section 7). Similarly to SDGLMC, they are all based on the assumption that the outcome follows a Poisson distribution. The main difference lies in how the exposure effect is modeled, and how unmeasured confounding is accounted for.
The linear predictor is a function of exposure and measured confounders. To control for unmeasured confounding, additional covariates and basis functions are included in the linear predictor (see, e.g., Peng et al., 2006, Peng and Dominici, 2008). Several approaches from the literature are then fitted under the Bayesian paradigm: (i) the Null model, which does not account for unmeasured confounding; (ii) the GLMadj model, which adjusts for unmeasured confounding through a natural cubic spline of time (see, e.g., Peng and Dominici, 2008, Zhou et al., 2021, Masselot et al., 2022); (iii) the JDZ model by Janes et al. (2007), which is similar to the previous approach but further decomposes the exposure effect in two orthogonal parts; (iv) the Dummy model (see, e.g, Dai et al., 2014, Klompmaker et al., 2023), which allows for season-specific exposure effects; and (v) the Periodic model by Peng et al. (2005) and Chen et al. (2013), which assumes that the exposure effect is a linear combination of trigonometric functions. A hierarchical structure is assumed for each of them, with spatial-specific coefficient vectors being independently and identically distributed. Further details are provided in Section D of the Supplementary Material.
6 Analysis of Artificial Data
To validate the proposed approach, we conduct two simulation studies, both inspired by the real data example presented in Section 2. In the first one, which is described next, we want to assess the ability of SDGLMC to mitigate the confounding bias.
Simulation Study 1: Confounding Bias Mitigation
We consider the ${N}=117$ contiguous HDs of Piemonte and Lombardia, then we assume that ${T}=600$ and there are no measured confounders (i.e., ${p_{m}}=0$). The exposure $X_{it}$ and the unmeasured confounder $Z_{it}$ are correlated vector autoregressive processes of order $1$ (VAR($1$)) with a linear correlation coefficient of 0.5, and with spatially-structured, temporally independent innovations (Cressie and Wikle, 2015, Sahu, 2022). A proper conditional autoregressive (PCAR) specification (Banerjee et al., 2014, Besag, 1974, Sahu, 2022) is assumed to model the spatial dependence of these innovations. The marginal spatio-temporal means of exposure and unmeasured confounder are generated from Fourier basis expansions. The correlation matrix of $X_{it}$ is parameterized by spatial and temporal autocorrelation parameters, i.e. $\phi_{x}^{S}\in(0,1)$ and $\phi_{x}^{T}\in(0,1)$, respectively. Similarly, $\phi_{z}^{S}$ and $\phi_{z}^{T}$ represent the autocorrelation parameters of $Z_{it}$, which is sampled conditional on $X_{it}$. These four parameters represent the strength of the spatial and temporal dependence, and define three scenarios that are summarized in Table 1.
It is worth noting that, in the concurvity setting, $X_{it}$ and $Z_{it}$ are considered as fixed unknown functions, particularly when exposure and unmeasured confounder vary at large scales. The stochastic generative model we use continues to be valuable within this framework of fixed functions because the realizations of $X_{it}$ and $Z_{it}$ provide reasonable values for exposure and unmeasured confounder that could arise in real-world applications where a reasonable stochastic process is absent (Paciorek, 2010).
The outcome variable is sampled from the Poisson distribution in Equation (1) using the following linear predictor:
${\color[rgb]{0,0,0}\vartheta_{it}=o_{it}+\beta_{0}+\beta_{1it}X_{it}+Z_{it}+u_{it},\quad u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})\,,}$
where $\tau_{u}^{2}=10^{-3}$. Additional information on the data generation mechanism is provided in Section E of the Supplementary Material.
Table 1: Simulation scenarios and corresponding definitions.
Scenario Unmeasured confounder, $Z_{it}$ $\phi_{x}^{S}$ $\phi_{z}^{S}$ $\phi_{x}^{T}$ $\phi_{z}^{T}$
S1 Smoother than exposure in space and time 0.2 0.98 0.2 0.98
S2 Smoother than exposure only in space 0.2 0.98 0.98 0.2
S3 Smoother than exposure only in time 0.98 0.2 0.2 0.98
The proposed approach is compared to the Null model and to the GLMadj model. The results from the other competing approaches are not reported here, but they can be found in Section E of the Supplementary Material. A natural cubic spline of time with $6$ degrees of freedom is used to account for unmeasured confounding in GLMadj. Following Peng et al. (2006), the degrees of freedom for the spline are chosen by fitting several models with different numbers of bases and selecting the minimum number from which the estimates appear stabilized.
Figure 3 depicts the results in terms of percent change of risk associated with a 10-unit increase in the exposure, specifically focusing on its constant and temporal components, i.e. it is defined as $P_{t}=100(\exp\{10(\overline{\delta}_{1}+\delta_{1t})\}-1)$. For each scenario, the black solid line represents the space-averaged true exposure effect. The estimate of $P_{t}$ obtained from SDGLMC is shown in red, with the solid line indicating its average, and the shaded area indicating the percentiles $2.5$ and $97.5$ of its distribution over 100 replicates.
The dashed gray and dashed green lines represent the pooled estimates from the Null and GLMadj models, averaged over the 100 replicates. Whereas the Null model is evidently biased in all cases, GLMadj and SDGLMC significantly reduce the confounding bias in Scenario S1, although without entirely removing it. In Scenario S2, the confounding bias reduction is still achieved by SDGLMC compared to the Null model, but to a lesser extent. This could be because the unmeasured confounder varies at high frequencies in time, so it cannot be captured by the time-varying intercepts that only account for gradual changes. Also, the difference $X_{it}-\widehat{X}_{it}$ is approximately as rough as $Z_{it}$, and, due to their correlation, a portion of the variability of the unmeasured confounder is explained by the local exposure effect. Scenario S3 is very similar to Scenario S1, suggesting that the lack of smoothness in space of the unmeasured confounder does not seem to alter the performance (in terms of confounding bias) of the proposed approach.
Figure 3: Simulation Study 1: space-averaged percent change of risk associated with a 10-unit increase in the exposure, $P_{t}$, for three scenarios and for the Null (dashed gray line), GLMadj (dashed green line), and SDGLMC (red solid line) models. The shaded area represents the percentiles $2.5$ and $97.5$ of the posterior means of $P_{t}$ obtained from SDGLMC over 100 replicated datasets. The black solid line represents the true exposure effect.
With regard to the variability of the exposure effect over space, we have computed the posterior means of the spatial effects, $\tilde{\bm{\delta}}_{1}$, for each scenario. The average of these posterior means over 100 replicated datasets are shown in Figures 46 and are compared to the (averaged) posterior means of the differences between area-specific and pooled estimates – i.e., the same quantity depicted in Figure 9 – obtained from the competing models.
All methods recover the simulated spatial pattern, with lower values in the middle of the study area. SDGLMC provides accurate estimates in all scenarios, while the Null and GLMadj models tend to underestimate the spatial variability in some cases.
Figure 4: Simulation Study 1: area-specific deviations from the time-varying percent change of risk associated with a 10-unit increase in the exposure, under Scenario S1.
Figure 5: Simulation Study 1: area-specific deviations from the time-varying percent change of risk associated with a 10-unit increase in the exposure, under Scenario S2.
Figure 6: Simulation Study 1: area-specific deviations from the time-varying percent change of risk associated with a 10-unit increase in the exposure, under Scenario S3.
The results above are confirmed also by Tables E.4–E.6 in the Supplementary Material, where the average mean squared error ($MSE_{avg}$), the maximum absolute bias ($MAB$), and the average mean absolute error ($MAE_{avg}$) are reported for each model and each scenario (see the Supplementary Material, Section E.2, for definitions). We also report these metrics for each component of the exposure effect to assess their contribution to the overall error. In general, we note that SDGLMC outperforms the competing models in all scenarios.
Simulation Study 2: Recovery of the Local Exposure Effect ($f^{\prime}$) and Model Reparameterization
In the second simulation study, described in Section F of the Supplementary Material, we want to assess whether the proposed approach is able to recover the first derivative, $f^{\prime}(\cdot;\bm{\gamma}_{f})$, when the functional form is assumed known. The data are generated using Equations (1)–(2), and we show that SDGLMC is able to capture the dynamics of the exposure effect, but there remains a residual bias which is due to confounding.
Another objective of this second simulation study is to motivate, from a practical viewpoint, the decomposition of the exposure effect into local-scale and large-scale components described in Equation (3). We compare SDGLMC with a reparameterized model that does not remove the overall trend from the exposure. As expected from the discussion in Section 3, the results suggest that the proposed approach is more effective in reducing the confounding bias and always has lower $MSE_{avg}$, $MAB$ and $MAE_{avg}$ values.
7 Application to the Italian Mortality Data
We apply SDGLMC to the Italian mortality dataset (see Section 2), and compare its results to those from the competing approaches. The large-scale trend, $\widehat{X}_{it}$, is obtained using a principal spline with $5$ spatial bases and $20$ temporal bases. For each HD and time instant, ${\bf M}_{it}$ contains bases from a natural cubic spline of daily temperature and daily relative humidity (with 7 and 4 degrees of freedom, respectively), and indicator variables for weekdays, holidays and lockdown periods (hence ${p_{m}}=19$; see also Section A of the Supplementary Material).
Types of Space-Time Interaction Component, $\delta_{0it}^{*}$
We can envision several distributional specifications for the space-time interaction component of the intercept, $\delta_{0it}^{*}$, depending on the correlation structure we want to impose. Knorr-Held (2000) proposes four types of models, giving rise to four types of space-time interaction terms, characterized by different prior assumptions. These are reviewed by Franco-Villoria et al. (2022) and Orozco-Acosta et al. (2023), and are summarized in Table 2. Type T1 assumes unstructured space-time variability, and is obtained by setting $\delta_{0it}^{*}=0$ for all $i$ and $t$, since an unstructured error is already present in Equation (7). Type T2 assumes independent temporal trends for each HD, with a common variance for the innovations. Type T3 assumes independent spatial trends for each time instant, with a common conditional variance. Type T4 assumes complete dependence, i.e. the temporal trends are different from HD to HD, but are more likely to be similar if the HDs are neighbors.
The prior assumptions discussed in Knorr-Held (2000) can be further relaxed. In fact, during the pandemic, some HDs in Lombardy – more specifically in the province of Bergamo – experienced a much greater increase in deaths than any other HD. Therefore, a fifth interaction type, T5, is also considered, assuming independent space-time interactions across the HDs that follow random walk processes with area-specific evolution variances. This captures temporal trends that differ considerably among HDs but do not have any spatial structure.
Table 2: Space-time interaction types considered for $\delta_{0it}^{*}$. DIC values for the corresponding models are reported, with the lowest value highlighted in italics.
Type Space-Time structure Distributional Assumption DIC
T1 Absent $\delta_{0it}^{*}=0$ 814038
T2 Independent RW, same variance $\delta_{0it}^{*}\sim\mathcal{N}(\delta_{0i,t-1}^{*},\sigma_{w0}^{*2})$ 807711
T3 Independent ICAR $\bm{\delta}_{0t}^{*}\sim ICAR(\sigma_{w0}^{*2})$ 811293
T4 Complete dependence $\bm{\delta}_{0t}^{*}-\bm{\delta}_{0,t-1}^{*}\sim ICAR(\sigma_{w0}^{*2})$ 811923
T5 Independent RW, different variances $\delta_{0it}^{*}\sim\mathcal{N}(\delta_{0i,t-1}^{*},\sigma_{w0i}^{*2})$ 803861
Model Fitting and Presentation of Results
The MCMC algorithm described in Section 4 is run for 100,000 iterations, and posterior inference is based on the final 80,000 draws. The model’s chains are monitored for convergence visually through trace plots and using the R-statistic of Gelman (1996) on two chains that are started from different initial points simultaneously.
After fitting a model for each of the five types of space-time interaction components, the one with the lowest deviance information criterion (DIC; Spiegelhalter et al., 2002) is selected. Based on the last column of Table 2, the model with interaction type T5 has the smallest DIC value, hence its results will be discussed below. A reason for this may be that, in comparison to the models proposed by Knorr-Held (2000), the model with T5 seems more able to capture very specific spatio-temporal trends, such as those observed in the Bergamo area due to the pandemic.
We assess the adequacy of the model fit by analyzing Bayesian $p$-values for each HD and time instant. The chosen measure evaluates how well the model aligns with the observed data, and involves calculating the quantity $\Pr\bigl{(}Y_{it}>Y^{rep}_{it}\bigr{)}+\frac{1}{2}\Pr\bigl{(}Y_{it}=Y^{rep}_{it}\bigr{)}$, where $Y^{rep}_{it}$ is generated from the predictive distribution in each MCMC iteration. Bayesian $p$-values close to $0.5$ suggest that the generated data align well with the model. Conversely, values approaching the boundaries of the unit interval, i.e. $<0.1$ or $>0.9$, indicate a poor fit (Gelman et al., 2013). The estimated mean $p$-value is around $0.5$ suggesting an adequate fit.
Figure 7 depicts the estimated space-averaged local exposure effect (i.e., $\overline{\delta}_{1}+\delta_{1t}$), expressed in terms of percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. The red solid line represents its posterior mean, while the shaded area the corresponding 95% credible interval (CI). The exposure effect is not constant in time, as no horizontal straight line can be drawn within the CI. This effect, instead, exhibits varying peaks across different years, and three phases are clearly identifiable: before (2018-19), during (2020-21) and after (2022) the pandemic outbreak.
During the first two years, the local exposure effect exhibits a seasonal pattern, with the peaks occurring during the summer months. The estimated association is bounded away from zero from August to September 2018 from April to October 2019, with the highest change in mortality risk of $5.50\%$ (95% CI: $[3.03\%,7.98\%]$) occurring on June 29, 2019.
Figure 7: Time-varying percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations estimated by SDGLMC. The red solid line represents its posterior mean, while the shaded area the corresponding 95% CI.
It is worth noting that a higher percent change of mortality risk occurs when the fine particulate matter is generally less present in the air, so this may indicate that a threshold below which the exposure effect is null might not exist (see, among others, Brunekreef et al., 2021b, Cork et al., 2023, Dominici et al., 2022, Renzi et al., 2021). The seasonal behavior seems to be confirmed by other studies (Nawrot et al., 2007, Peng et al., 2006, Roberts, 2004, Stafoggia et al., 2008, Sun et al., 2015, Tian et al., 2018).
These seasonal changes of the risk suggest that the unmeasured confounders could have not only a marginal effect on mortality (hence influencing $\beta_{0it}$), but also an interaction effect with the exposure (hence affecting $\beta_{1it}$). In fact, the seasonal pattern might be due to potential interactions between PM2.5 concentrations and air temperature. As noted in Section 1, increased ventilation and physiological susceptibility to air pollution may explain the higher percent change of risk during summer. Besides, interaction with other variables, such as other pollutants, are also possible (see, e.g., Liu et al., 2023, Ji et al., 2022, Siddika et al., 2019).
Previous studies have also reported that particle composition may modify the effect of PM2.5 on mortality. In fact, fine particulate matter is a complex mixture of organic (such as black carbon and organic carbon) and inorganic compounds (such as sulfate, nitrate, and ammonium), and the relative proportion of these components may vary in time (Franklin et al., 2008, Masselot et al., 2022, Vinson et al., 2020). Indeed, their precursors are emitted from both natural (e.g., oceans and volcanoes) and anthropogenic sources, such as fossil fuel combustion, industrial processes, and residential heating, and their transportation and transformation to PM2.5 in the atmosphere are mainly influenced by meteorological conditions (Hao et al., 2023, Jing et al., 2020, Kelly and Fussell, 2012). Toxicological studies indicate that the PM constituents have different effects on human health (see, e.g., Kelly and Fussell, 2012), and some authors have suggested that ammonium is probably among the most harmful components (Masselot et al., 2022, Hao et al., 2023).
During the pandemic phase, the CI always contains the zero except for few days during December 2021. The literature reveals a lack of consensus regarding the association between air pollution and adverse COVID-19 health outcomes in the general population (see, for example, Brunekreef et al., 2021a, Walton et al., 2021). More recently, the review by Bhaskar et al. (2023) suggests a positive association in the majority of previous studies. Our findings, instead, indicate that the local exposure effect in 2020 and 2021 is generally not bounded away from zero (except for few days). To justify this, consider first of all that our model uses all-cause mortality in the elderly population as the health outcome, so it is not directly comparable to the COVID-19 studies mentioned earlier. Secondly, the issuance of the lockdown measures by the Italian government, as well as the mandatory use of facial masks, not only effectively curbed the transmission of the SARS-CoV-2 virus (and of other viruses), but also potentially reduced exposure to air pollution. During the lockdown periods, people were required to stay at home, resulting in reduced exposure to outdoor air pollutants. The ventilation was limited because the lockdowns occurred mainly during winter months. Moreover, the use of facial masks may have reduced the inhalation of airborne hazardous particles (Kodros et al., 2021, Zhou et al., 2018).
Although lockdown periods were not issued in the second half of 2021, the local exposure effect remains not bounded away from zero. This could be due to the fact that people continued to adopt measures such as wearing masks and sanitizing hands. Additionally, it is possible that the phenomenon of mortality displacement or harvesting may have occurred. This is a consequence of the fact that the most vulnerable people may have died earlier than expected in the preceding years due to the SARS-CoV-2 virus in 2020-2021, resulting in an apparently null (or even protective) effect of the exposure afterwards (Armstrong et al., 2017, Bhaskaran et al., 2013, Zanobetti and Schwartz, 2008, Zanobetti et al., 2000).
During the third phase, people started to return to their normal activities, resulting in increased exposure to air pollution. In fact, the local exposure effect has a shape similar to that in the first phase, and is bounded away from zero from July to September 2022, with the highest change in mortality risk of $3.53\%$ (95% CI: $[0.61\%,6.44\%]$) occurring on July 26.
Comparison with Competing Approaches
For each approach introduced in Section 5, Figure 8 shows the pooled estimated percent change of risk associated with a $10\ \mu g/m^{3}$ increase in PM2.5 levels. The posterior mean from SDGLMC is also shown (in red) for comparison. The Null model yields the highest estimate on average ($3.92\%$), whereas the GLMadj model estimates the lowest percent change on average ($0.21\%$). The models (i)–(iii) assume that the association of interest is constant over time, however we have argued that this assumption is unlikely.
Figure 8: Estimated percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations: comparison of SDGLMC with the approaches described in Section 5.
The only two approaches able to capture time-varying strengths of the association are the Dummy and the Periodic models. They seem to agree with SDGLMC on the temporal evolution of the effect, showing peaks in summer and troughs in winter. The Dummy and the Periodic models, however, have some limitations, since both assume that the effect changes within the year, but is the same across years. This is not consistent with the evidence from the data, as the pandemic outbreak may have altered the strength of the association.
Although the Dummy model allows for abrupt changes in the exposure effect, these are unlikely in practice as pointed out by Bhaskaran et al. (2013) and Peng et al. (2005). Moreover, the results may be too sensitive to the choice of the boundaries of each season, and a definition based on the calendar seasons may result suboptimal.
In terms of spatial variation, we analyze the differences between area-specific and pooled estimates of the percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. Figure 9 shows the posterior means of those differences under each competing approach. It also shows the posterior mean of the spatial component of the local exposure effect, i.e. $\delta_{1i}$, expressed in terms of percent change of risk. The spatial variability of the exposure effect approaches zero in all cases except for the Null and GLMadj models. However, when considering the associated uncertainty, there is no evidence of spatial variation in the exposure effect for any of the models. This may be related to the relatively small size of the study region and the similarity between Piemonte and Lombardia (e.g., in terms of socio-economic indicators).
Figure 9: Spatial component of the percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations under each approach.
Comparison with a Restricted Model
In Section G of the Supplementary Material we want to understand the consequences of removing the measured confounders (i.e., temperature, relative humidity, and indicators for weekdays, national holidays, and lockdown periods) from the model specification. Therefore, we impose the restriction that the elements of $\bm{\alpha}$ in Equation (6) are zero. The results are insightful, as the estimated exposure effect greatly overlaps the estimate obtained when the measured confounders are included, with differences occurring only in summer. This suggests that SDGLMC can provide accurate estimates even in complex scenarios, and that the confounding effect may be time-varying and may be due to heat waves.
8 Discussion
Motivated by a real case-study, in this paper we have proposed the spatial dynamic generalized linear model for confounding (SDGLMC) as a solution for two types of bias that may affect the estimation and interpretation of the exposure–outcome relationship: misspecification and confounding bias. Misspecification bias arises from assuming that the exposure effect is constant in time when it is not. When analyzing the health effects of air pollution, Chiogna and Gaetan (2002) and Peng et al. (2005), among others, argue that there is uncertainty about the “true” functional form of the exposure–outcome association. Therefore, assuming a constant effect may be too restrictive and result in biased estimates. We acknowledge this fact and assume that there exists a smooth (but unknown) function of the exposure, $f(\cdot;\cdot)$, describing the association of interest. Confounding bias is due to concurvity. To control for it, we decompose the exposure effect into large- and small-scale components, then we want to recover the local exposure effect. The same objective is pursued by Janes et al. (2007) and Dupont et al. (2022), but, differently from these authors, we assume that the local effect is spacetime-varying. Hence, SDGLMC constitutes an extension of their models. Importantly, another connection exists between SDGLMC and the work of Dupont et al. (2022), as the smooth function of spatial coordinates in their model and the spatio-temporal intercepts in SDGLMC play a similar role.
We have shown that SDGLMC enjoys some advantages. On one hand, it is able to capture any spacetime-varying exposure effects, thus overcoming the misspecification bias. On the other hand, thanks to the first-order Taylor’s expansion, SDGLMC can model the local exposure effect within a flexible framework. To recover the local, short-term association, the outcome is regressed on a detrended exposure. This is useful because detrending improves estimation stability and helps reduce the confounding bias (Janes et al., 2007, Greven et al., 2011).
The effectiveness of the proposed approach has been demonstrated through artificial data analyses. We have shown that it performs better than several competitors. In particular, SDGLMC not only can reduce the bias, but also can capture complex spatio-temporal dynamics of the exposure effect and almost recover its true functional form. The bias is less alleviated when the unmeasured variable is less smooth than the exposure in time, but it is worth noticing that this situation is rather unrealistic in practice. Note that the models assuming a constant association (e.g., GLMadj) could achieve better performances if the true effect is constant, however the shape of the exposure–outcome association is typically unknown. Unless there is strong prior information about this shape, we thus suggest using SDGLMC.
Many previous empirical works assume a constant exposure–outcome association. However, the evidence from the analysis of the Italian mortality data have indicated that this assumption is conceivably incorrect. In fact, the estimated local exposure effect exhibits a seasonal pattern, with a higher percent change of risk observed during the summer months, when the overall PM2.5 level is lower. Similar conclusions were reached in other studies (Peng et al., 2005, 2006, Roberts, 2004, Stafoggia et al., 2008, Sun et al., 2015, Tian et al., 2018). Although this fact is apparently counterintuitive, it has been observed that it might be due to possible interaction effects of PM2.5 concentrations with meteorological variables, such as temperature, and other air pollutants. We have also noted that our result may be related to possible effect modification due to changes in the particulate matter mixture (Franklin et al., 2008, Hao et al., 2023, Masselot et al., 2022). The relative proportions of the PM2.5 components have not been analyzed in this study, so they could be the subject of future research. However, it is important to note that Hao et al. (2023) found a non-linear long-term association between some PM2.5 constituents and mortality.
A reviewer pointed out that if the effect modification is solely due to measured confounders, it can be studied more straightforwardly by including interaction terms between PM2.5 and the confounders in the linear predictor. We explore this possibility in Section G.2 of the Supplementary Material. However, since the estimated effect is significantly different from that in Figure 7 (especially during the summer months and the pandemic), we conclude that the effect modification is due to both measured and unmeasured confounders.
An exception to the seasonal behavior are the years 2020 and 2021, when the exposure effect is not bounded away from zero. This can be justified by the fact that the lockdown measures, such as stay-at-home orders and the use of masks, have likely reduced the exposure to air pollution. A similar result is found in the absence of measured confounders (see Figure G.1 in the Supplementary Material), and is also suggested by recent studies (see, e.g., Hadei et al., 2021, López-Feldman et al., 2021). It is worth recalling that none of the competing approaches in Section 2 were able to capture this behavior. However, separate analyses per year (see Table 3) suggest patterns that are qualitatively similar to our results. We limited this analysis to the GLMadj and JDZ models due to the limited temporal extent. Nevertheless, both models indicate a non-significant exposure effect during the second phase.
Table 3: Posterior means obtained from fitting the GLMadj and JDZ approaches separately for each year. The values are expressed in terms of percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. The 95% CIs are reported within the square brackets.
2018 2019 2020 2021 2022
GLMadj 0.75 [0.06, 1.45] 0.81 [0.22, 1.39] 0.14 [-0.43, 0.69] -0.03 [-0.69, 0.68] 0.67 [0.01, 1.31]
JDZ 0.82 [0.13, 1.50] 0.85 [0.25, 1.43] 0.26 [-0.30, 0.81] -0.03 [-0.69, 0.68] 0.58 [-0.07, 1.22]
To account for the residual spatial autocorrelation in the outcome, possibly due to unmeasured confounding, SDGLMC accounts for the adjacency of areal units. That is, spatial and space-time components of the intercept are included into the linear predictor. Besides, SDGLMC allows to possibly capture the spatial variations of the local exposure effect, which could potentially discover differences due to urbanicity and morphology of the study area, for example. However, in the application to the Italian mortality data, the local exposure effect is homogeneous across space, probably because the study area is relatively small, and Piemonte and Lombardia have similar socio-economic and environmental characteristics.
The decomposition of the exposure effect into two orthogonal components allows Janes et al. (2007) to test for the presence of unmeasured confounders. The authors state that if the overall and local exposure effects are equal, then it is likely that the confounding bias is not present. In our case, this testing procedure is theoretically possible, but it is not straightforward to implement in practice, since specific assumptions on the functional form of $f(\cdot;\bm{\gamma}_{f})$ are required.
It is arguable that, even if SDGLMC could recover the true exposure effect, $f(\cdot;\cdot)$, its functional form still remains unknown. In other words, the complete list of interaction terms cannot be determined. However, we believe that this is not a drawback only of the proposed model. In fact, the list of interaction terms is unknown when fitting any regression model, and it is up to the researcher to choose the most appropriate form for $f(\cdot;\cdot)$. This decision may involve some degree of uncertainty. The proposed approach, instead, allows for more flexibility and should be able to capture the effect of lockdowns as well as interaction effects between the exposure and other variables, such as other pollutants, so that inferential problems due to misspecification are overcome.
Further research could focus on the causal interpretation of the estimated exposure effect, in which case other assumptions are needed (see, e.g., Hernán and Robins, 2020). However, note that the analytical approach pursued here should not be considered as less credible than causal inference methods (Dominici and Zigler, 2017).
In conclusion, SDGLMC is a flexible and effective approach to estimate the exposure–outcome association in the presence of spacetime-varying effects and unmeasured spatio-temporal confounding. The model is able to capture the complex dynamics of the exposure effect, and to control for confounding. The proposed approach is a valuable tool for researchers interested in the health effects of air pollution, and it can be extended to other fields where the exposure–outcome association is of interest.
Acknowledgements
Hersbach et al. (2018) was downloaded from the C3S (2023). INERIS et al. (2022) was downloaded from the CAMS (2023). The results contain modified C3S and CAMS information. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains. The research of C.Z., P.V., and L.I. is funded by the European Union - NextGenerationEU. C.Z.: PRIN PNRR 2022 (Bando 2023) SLIDE - Stochastic Modeling of Compound Events (P2022KZJTZ - CUP D53D23018920001). P.V. and L.I.: research project PRIN2022 CoEnv - Complex Environmental Data and Modeling (2022E3RY23 - CUP D53D23011080006). A.M.S. is grateful for financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Discovery Grant RGPIN-2017-04999).
Supplementary Materials
Appendix A Data Collection
We gather daily data to analyze the short-term association between PM2.5 and all-cause mortality in Piemonte and Lombardia regions for people aged 65 or more. The data span the period between 2018 and 2022, and are collected at the health district (HD) level.
  • Mortality data. All-cause mortality data are collected from the Italian National Institute of Statistics (ISTAT, 2023a) for each municipality in the two regions, and are then spatially aggregated to the HD level, which represents the areal unit. The spatial domain comprises ${N}=117$ HDs where $84$ belong to the Lombardia region and $33$ to the Piemonte region. The municipality of Campione d’Italia is excluded from the analysis since it has a small area and the time series contains many zeroes. Since the data refer to five consecutive years, the temporal dimension is ${T}=1826$.
    For $i=1,\dots,{N}$ and $t=1,\dots,{T}$, the standardized mortality ratio (SMR) is calculated as the ratio between the observed number of deaths, $Y_{it}$, and the expected number of deaths, $E_{it}$. The latter is obtained by indirect standardization (Banerjee et al., 2014, Sahu, 2022) as follows. The resident population aged 65 or more represents the population at risk and is obtained from the Italian National Institute of Statistics for each year (ISTAT, 2023c). The population is stratified by age class (65–69, 70–74, 75–79, 80–84, 85–89, 90–94, or 95+ years) and gender (male or female). For each stratum $k=1,\dots,K$, the mortality rate for the northwest Italy is retrieved from the Italian National Institute of Statistics for each year (ISTAT, 2023b). The expected counts are thus computed as follows:
    $E_{it}=\sum_{k=1}^{K}Pop_{itk}r_{tk}\,,$
    where $Pop_{itk}$ and $r_{tk}$ are the population at risk and the mortality rate in the $k$-th stratum, respectively.
  • Exposure data. Data on PM2.5 concentrations (expressed in $\mu g/m^{3}$) are collected from the Copernicus Atmosphere Monitoring Service’s (CAMS) Atmosphere Data Store (INERIS et al., 2022, CAMS, 2023). Hourly data are available at a spatial resolution of $0.1^{\circ}\times 0.1^{\circ}$, so we use zonal statistics to aggregate the estimates to the HD level, and then average across time to obtain daily data.
  • Potential confounders. Daily temperature, daily relative humidity, indicator variables for the days of the week, for national holidays and for lockdown periods are treated as potential measured confounders. We gather satellite measurements of air temperature and dew-point temperature (both expressed in Kelvin degrees) from the ERA5 reanalyses dataset available at the Copernicus Climate Change Service’s (C3S) Climate Data Store (C3S, 2023, Hersbach et al., 2018). Relative humidity ($RH$) is calculated from air temperature ($Temp$) dew-point temperature ($DT$) using the August-Roche-Magnus approximation formula (Alduchov and Eskridge, 1996):
    $RH=100\times\exp\left(\frac{17.625\times DT}{243.04+DT}-\frac{17.625\times Temp}{243.04+Temp}\right)\,.$
    These hourly data are spatially and temporally averaged, similarly to the exposure.
    The lockdowns are defined as follows: the first period is from March 9, 2020, to May 3, 2020, the second one is from December 21, 2020, to January 15, 2021, and the third one from March 6, 2021, to April 25, 2021.
Appendix B Descriptive Statistics
Descriptive statistics on the SMR data in all HDs are reported in Table B.1. The average SMR is $1.28$ and $1.18$ in Lombardia and Piemonte, respectively, and the standard deviation is slightly higher in Lombardia ($0.34$) than in Piemonte ($0.26$).
Table B.1: Descriptive statistics of daily SMR across all HDs in 2018–2022. Q1, Q2, and Q3 stand for first, second, and third quartiles, respectively.
Region Mean SD Min Q1 Q2 Q3 Max
Lombardia 1.28 0.29 0.75 1.14 1.25 1.39 3.52
Piemonte 1.18 0.22 0.68 1.03 1.14 1.29 2.10
Overall 1.25 0.26 0.75 1.12 1.22 1.37 3.09
Descriptive statistics on PM2.5 levels in all HDs during the whole study period are reported in Table B.2. The average daily PM2.5 concentration is $18.21\,\mu g/m^{3}$ (standard deviation = $10.36$) over the entire spatial domain, but Lombardia is more polluted than Piemonte. Table B.2 also reports descriptive statistics of temperature and relative humidity. The average daily temperature is $284.91\,K$ ($7.60$), while the average daily relative humidity is $91.76\%$ ($3.59$). These are very similar to the region-specific values (not shown).
Table B.2: Descriptive statistics of daily PM2.5 concentrations, temperature, and relative humidity across all HDs in 2018–2022. Q1, Q2, and Q3 stand for first, second, and third quartiles, respectively.
Variable Mean SD Min Q1 Q2 Q3 Max
PM2.5 ($\mu g/m^{3}$) 18.21 10.36 2.32 10.49 15.20 24.12 63.07
Temperature ($K$) 284.91 7.60 265.15 278.17 284.68 292.02 301.59
Relative humidity (%) 91.76 3.59 79.55 89.43 92.12 94.54 98.43
Finally, Table B.3 reports the annual death counts and annual averages of the daily SMR and daily PM2.5 for each administrative region.
Table B.3: Annual mortality counts, and annual averages of daily SMR and daily PM2.5 concentrations for Piemonte and Lombardia (2018–2022).
Year
Variable Region 2018 2019 2020 2021 2022
Mortality Counts Lombardia 89742 90613 123586 97365 101069
Piemonte 48739 48479 60364 51202 53581
Overall 138481 139092 183950 148567 154650
SMR Lombardia 1.32 1.32 1.23 1.29 1.26
Piemonte 1.20 1.20 1.16 1.18 1.17
Overall 1.29 1.29 1.21 1.26 1.24
PM2.5 ($\mu g/m^{3}$) Lombardia 18.56 20.41 18.36 19.42 20.30
Piemonte 14.00 16.47 14.33 15.17 15.91
Overall 17.27 19.30 17.22 18.22 19.06
Appendix C Posterior Full Conditional Distributions for the Proposed DGLM
We provide the computational details to make inference on the model proposed in Section 3 and following the discussion in Section 4 of the main manuscript. In particular, stacking Equation (7) of the main text over the ${T}$ time instants, we have (see also Chan and Jeliazkov, 2009):
$\bm{\vartheta}=\mathbf{o}+\mathbf{G}_{1}\bm{\gamma}_{1}+\mathbf{G}_{2}\bm{\gamma}_{2}+\mathbf{G}_{3}\bm{\gamma}_{3}+\mathbf{G}_{4}\bm{\gamma}_{4}+\tilde{\mathbf{u}}\,,$
where $\bm{\vartheta}=(\bm{\vartheta}_{1}^{\prime},\dots,\bm{\vartheta}_{T}^{\prime})^{\prime}$, $\mathbf{o}=(\mathbf{o}_{1}^{\prime},\dots,\mathbf{o}_{T}^{\prime})^{\prime}$, $\tilde{\mathbf{u}}=(\tilde{\mathbf{u}}_{1}^{\prime},\dots,\tilde{\mathbf{u}}_{T}^{\prime})^{\prime}$, $\bm{\gamma}_{1}=(\bm{\delta}_{01}^{*\prime},\bm{\delta}_{02}^{*\prime},\dots,\bm{\delta}_{0{T}}^{*\prime})^{\prime}$, $\bm{\gamma}_{2}=(\delta_{01},\delta_{11},\dots,\delta_{0{T}},\delta_{1{T}})^{\prime}$, $\bm{\gamma}_{3}=(\tilde{\bm{\delta}}_{0}^{\prime},\tilde{\bm{\delta}}_{1}^{\prime})^{\prime}$, $\bm{\gamma}_{4}=(\overline{\delta}_{0},\overline{\delta}_{1},\bm{\alpha}^{\prime})^{\prime}$. The observation matrices are $\mathbf{G}_{1}={\bf I}_{{N}{T}}$,
$\mathbf{G}_{2}=\begin{bmatrix}{\bf 1}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}1}&&&\\ &&{\bf 1}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}2}&&\\ &&&\ddots&\ddots&\\ &&&&{\bf 1}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}{T}}\end{bmatrix}\,,\qquad\mathbf{G}_{3}=\begin{bmatrix}{\bf I}_{N}&\mbox{diag}(\mathbf{\tilde{X}}_{\,\bm{\cdot}1})\\ \vdots&\vdots\\ {\bf I}_{N}&\mbox{diag}(\mathbf{\tilde{X}}_{\,\bm{\cdot}{T}})\end{bmatrix}\,,\qquad\mathbf{G}_{4}=\begin{bmatrix}{\bf 1}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}1}&{\mathbf{M}}_{1}\\ \vdots&\vdots&\vdots\\ {\bf 1}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}{T}}&{\mathbf{M}}_{T}\end{bmatrix}\,.$
Posterior samples are obtained using a Gibbs sampler with the following full conditionals, where the symbol $\bullet$” denotes data and all the parameters except for the parameter that is being updated.
  1. 1.
    Sample regression coefficients, $\bm{\gamma}_{j}$, for $j=1,2,3,4$:
    $\bm{\gamma}_{j}|\bullet\sim\mathcal{N}(\overline{\mathbf{V}}_{j}\overline{\bm{\mu}}_{j},\overline{\mathbf{V}}_{j})\,,$
    where $\overline{\mathbf{V}}_{j}=(\mathbf{K}_{j}+\mathbf{G}_{j}^{\prime}\mathbf{G}_{j}/\tilde{\tau}_{u}^{2})^{-1}$ and $\overline{\bm{\mu}}_{j}=\mathbf{G}_{j}^{\prime}\bm{\eta}_{j}/\tilde{\tau}_{u}^{2}$, where $\bm{\eta}_{j}=\bm{\vartheta}-\mathbf{o}-\sum_{l\neq j}\mathbf{G}_{l}\bm{\gamma}_{l}$, $\mathbf{K}_{j}=\mathbf{H}_{j}^{\prime}\mathbf{S}_{j}^{-1}\mathbf{H}_{j}$, and
    $\mathbf{H}_{1}=\begin{bmatrix}\mathbf{I}_{N}&&&\\ -\mathbf{I}_{N}&\mathbf{I}_{N}&\\ &\ddots&\ddots&\\ &&-\mathbf{I}_{N}&\mathbf{I}_{N}\\ \end{bmatrix}\quad\mbox{and}\quad\mathbf{S}_{1}^{-1}=\begin{bmatrix}\mathbf{V}_{\delta}^{*-1}&&&\\ &\sigma_{w0}^{*-2}(\mathbf{D-W})&&\\ &&\ddots&\\ &&&\sigma_{w0}^{*-2}(\mathbf{D-W})\\ \end{bmatrix}\,,$
    $\mathbf{H}_{2}=\begin{bmatrix}\mathbf{I}_{2}&&&\\ -\mathbf{I}_{2}&\mathbf{I}_{2}&\\ &\ddots&\ddots&\\ &&-\mathbf{I}_{2}&\mathbf{I}_{2}\end{bmatrix}\quad\mbox{and}\quad\mathbf{S}_{2}^{-1}=\begin{bmatrix}\mbox{diag}(V_{\delta 0}^{-1},V_{\delta 1}^{-1})&&&\\ &\mbox{diag}(\sigma_{w0}^{-2},\sigma_{w1}^{-2})&&\\ &&\ddots&\\ &&&\mbox{diag}(\sigma_{w0}^{-2},\sigma_{w1}^{-2})\\ \end{bmatrix}\,,$
    $\mathbf{H}_{3}={\bf I}_{2{N}}\quad\mbox{and}\quad\mathbf{S}_{3}^{-1}=\begin{bmatrix}\sigma_{\delta 0}^{-2}(\mathbf{D-W})&\\ &\sigma_{\delta 1}^{-2}(\mathbf{D-W})\\ \end{bmatrix}\,,$
    $\mathbf{H}_{4}={\bf I}_{{p_{m}}+2}\quad\mbox{and}\quad\mathbf{S}_{4}^{-1}=10^{-6}{\bf I}_{{p_{m}}+2}\,.$
  2. 2.
    Sample variance parameters of temporal effects, for $k=0,1$:
    $\sigma_{wk}^{2}|\bullet\sim\mathcal{IG}\left(a_{wk}+\frac{{T}-1}{2},b_{wk}+\frac{1}{2}\sum_{t=2}^{{T}}(\delta_{kt}-\delta_{k,t-1})^{2}\right)$
  3. 3.
    Sample variance parameters of spatial effects, for $k=0,1$:
    $\sigma_{\delta k}^{2}|\bullet\sim\mathcal{IG}\left(a_{\delta k}+\frac{{N}-1}{2},b_{\delta k}+\frac{1}{2}\tilde{\bm{\delta}}_{k}^{\prime}(\mathbf{D}-\mathbf{W})\tilde{\bm{\delta}}_{k}\right)$
  4. 4.
    Sample variance parameter of spatio-temporal effect:
    $\sigma_{w0}^{*2}|\bullet\sim\mathcal{IG}\left(a_{w0}^{*}+\frac{({N}-1)({T}-1)}{2},b_{w0}^{*}+\frac{1}{2}\sum_{t=2}^{{T}}(\bm{\delta}_{0t}^{*}-\bm{\delta}_{0,t-1}^{*})^{\prime}(\mathbf{D}-\mathbf{W})(\bm{\delta}_{0t}^{*}-\bm{\delta}_{0,t-1}^{*})\right)$
  5. 5.
    Sample the latent variables, $\vartheta_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$, using Metropolis-Hastings steps with a Gaussian proposal distribution:
    $\vartheta_{it}|\bullet\propto\exp\left\{Y_{it}\vartheta_{it}-\exp(\vartheta_{it})-\frac{1}{2\tilde{\tau}_{u}^{2}}(\vartheta_{it}-o_{it}-\vartheta_{it}^{*})^{2}\right\}\,,$
    where $\vartheta_{it}^{*}$ is a generic element of $\bm{\vartheta}_{t}^{*}=(\vartheta_{11}^{*},\vartheta_{21}^{*},\dots,\vartheta_{{N}{T}}^{*})^{\prime}={\bf 1}_{N}(\overline{\delta}_{0}+\delta_{0t})+\tilde{\bm{\delta}}_{0}+\bm{\delta}_{0t}^{*}+\mathbf{\tilde{X}}_{t}(\overline{\delta}_{1}+\delta_{1t})+\mathbf{\tilde{X}}_{t}\odot\tilde{\bm{\delta}}_{1}+{\mathbf{M}}_{t}\bm{\alpha}$.
  6. 6.
    Sample the measurement error variance:
    $\tilde{\tau}_{u}^{2}|\bullet\sim\mathcal{IG}\left(a_{u}+\frac{{N}{T}}{2},b_{u}+\frac{1}{2}\sum_{t=1}^{{T}}(\bm{\vartheta}_{t}-\mathbf{o}_{t}-\bm{\vartheta}_{t}^{*})^{\prime}(\bm{\vartheta}_{t}-\mathbf{o}_{t}-\bm{\vartheta}_{t}^{*})\right)\,.$
Appendix D Details on Competing Models
The proposed DGLM is compared to several approaches from the literature. These are fitted under the Bayesian paradigm, and the general hierarchical structure is as follows, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$:
$Y_{it}|\vartheta_{it}$ $\sim\mathcal{P}oi(\exp(\vartheta_{it}))$
$\vartheta_{it}$ $=o_{it}+\mathbf{A}_{it}^{\prime}\mathbf{a}_{i}+\mathbf{C}_{it}^{\prime}\mathbf{c}_{i}+\varepsilon_{it},\quad\varepsilon_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{\varepsilon}^{2})$
$\mathbf{a}_{i}$ $\sim\mathcal{N}(\bm{\mu}_{a},\bm{\varSigma}_{a})$
$\bm{\mu}_{a}$ $\sim\mathcal{N}(0,10^{6}{\bf I}_{p_{a}})$
$\bm{\varSigma}_{a}$ $\sim\mathcal{IW}(p_{a}+1.1,10^{-4}{\bf I}_{p_{a}})$
$\mathbf{c}_{i}$ $\sim\mathcal{N}({\bf 0},10^{6}{\bf I}_{p_{c}})$
$\tau_{\varepsilon}^{2}$ $\sim\mathcal{IG}(1,2.1\times 10^{-5})$
where $\mathbf{A}_{it}$ is a $p_{a}$-dimensional vector of covariates related to the exposure effect, $\mathbf{C}_{it}$ is a $p_{c}$-dimensional vector of covariates (i.e., intercept and measured confounders), and $\mathbf{a}_{i}$ and $\mathbf{c}_{i}$ are the corresponding vectors of coefficients. The vector $\bm{\mu}_{a}$ and the matrix $\bm{\varSigma}_{a}$ are of main interest since they correspond to the overall exposure effect and the between-site variability, respectively.
The specification of $\mathbf{A}_{it}$ and $\mathbf{C}_{it}$ depends on the approach under consideration as listed below.
  • Null: the model that does not account for unmeasured confounding, for which $\mathbf{A}_{it}=X_{it}$ and $\mathbf{C}_{it}=(1,{{\bf M}}_{it}^{\prime})^{\prime}$. The regression coefficients are allowed to vary across areas but not in time.
  • GLMadj: the model that includes smooth function of time, $h(t;df)$, modelled as a natural cubic spline with $df$ degrees of freedom. Let $\mathbf{B}_{t}$ denote the vector of basis functions evaluated at time $t$. Then, we define $\mathbf{A}_{it}=X_{it}$ and $\mathbf{C}_{it}=(1,{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$.
  • Dummy: Dai et al. (2014), Klompmaker et al. (2023) assume that the long-term exposure–outcome association vary by season. Following also Chen et al. (2013), Peng et al. (2005), we may extend this approach to our setting by including interaction terms. Let $D_{jt}$ indicate an indicator variable for the $j$-th calendar season, for $j=1,\dots,4$, that is:
    $D_{jt}=\begin{cases}1&\mbox{if the day $t$ is in the $j$-th season}\\ 0&\mbox{otherwise}\end{cases}$
    Then, $\mathbf{A}_{it}=(X_{it}D_{1t},\dots,X_{it}D_{4t})^{\prime}$ and $\mathbf{C}_{it}=(1,{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$. Henceforth, the Dummy model is similar to GLMadj, but it also captures season-specific exposure effects.
  • Periodic: the exposure effect is assumed to be a linear combination of trigonometric functions, as in Peng et al. (2005), Chen et al. (2013). In this case, $\mathbf{A}_{it}=(X_{it},X_{it}\cos(2\pi t/365),X_{it}\sin(2\pi t/365))^{\prime}$ and $\mathbf{C}_{it}=(1,{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$.
  • JDZ: the model proposed by Janes et al. (2007), where the local exposure effect is of interest, hence $\mathbf{A}_{it}=(X_{it}-\widehat{X}_{it})^{\prime}$ and $\mathbf{C}_{it}=(1,\widehat{X}_{it},{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{*\prime})^{\prime}$. Here, $\widehat{X}_{it}$ denotes the overall temporal trend (see also Section 3), and $\mathbf{B}_{t}^{*}$ denotes the vector of basis functions from a natural cubic spline with $df-1$ degrees of freedom, evaluated at time $t$.
The unknown function $h(\cdot;\cdot)$ is modelled as a natural cubic spline of time with $df=43$ degrees of freedom (i.e., 8 equally-spaced knots per year plus 3 more knots to account for the peaks in the SMR time series), and is used to control for unmeasured confounding. The overall trend, $\widehat{X}_{t}$, is obtained as the fitted values of a regression of $X_{it}$ on a natural cubic spline of time with $4$ degrees of freedom per year. In the real data application, bases from natural cubic splines of temperature (with 7 degrees of freedom) and of relative humidity (with 4 degrees of freedom) are included as measured confounders. ${{\bf M}}_{it}$ also includes the indicator variables for the day of the week, for holidays, and for the lockdown periods.
Appendix E Additional Details on Simulation Study 1
E.1 Data Generation Mechanism
To simulate the variables, we use the following data generation mechanism. The spatial domain comprises ${N}=117$ contiguous HDs, while the temporal dimension is set to ${T}=600$. Let ${\bf X}=({\bf X}_{1},\dots,{\bf X}_{{T}})^{\prime}$, with ${\bf X}_{t}=(X_{1t},\dots,X_{{N}t})^{\prime}$, $t=1,\dots,{T}$, be the exposure, and let ${\bf Z}=({\bf Z}_{1},\dots,{\bf Z}_{{T}})^{\prime}$, with ${\bf Z}_{t}=(Z_{1t},\dots,Z_{{N}t})^{\prime}$, $t=1,\dots,{T}$, be the unmeasured confounder. It is thus assumed that
$\begin{bmatrix}\mathbf{X}\\ \mathbf{Z}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\bm{\mu}_{x}\\ \bm{\mu}_{z}\end{bmatrix},\begin{bmatrix}\bm{\varSigma}_{x}&\bm{\varSigma}_{xz}\\ \bm{\varSigma}_{zx}&\bm{\varSigma}_{z}\end{bmatrix}\right)\,,$ (11)
where $\bm{\mu}_{x}$ and $\bm{\mu}_{z}$ are the spatio-temporal means generated from Fourier basis expansions. It is assumed that the covariance matrices can be factorized as $\bm{\varSigma}_{x}=\bm{\varSigma}_{x}^{1/2}\bm{\varSigma}_{x}^{1/2\prime}$, $\bm{\varSigma}_{z}=\bm{\varSigma}_{z}^{1/2}\bm{\varSigma}_{z}^{1/2\prime}$, $\bm{\varSigma}_{xz}=\bm{\varSigma}_{zx}^{\prime}=\rho_{xz}\bm{\varSigma}_{x}^{1/2}\bm{\varSigma}_{z}^{1/2\prime}$, and $\rho_{xz}={\rm\mathbb{C}or}(\mathbf{X},\mathbf{Z})$.
We then assume that exposure and unmeasured confounder are spatial VAR(1) processes (Cressie and Wikle, 2015, Sahu, 2022). We now explain how $\bm{\varSigma}_{x}$ is parameterized, and the same applies to $\bm{\varSigma}_{z}$. If a stable VAR(1) model is assumed for the exposure where the temporal autocorrelation is controlled by a scalar, $\phi_{x}^{T}\in(0,1)$, then it can be shown (see, for example, Lütkepohl, 2005, Section 2.1.4) that $\bm{\varSigma}_{x}=\bm{\Gamma}_{x,ar}\otimes{\bf C}_{x}(0)$, where
${\bf C}_{x}(0)={\rm\mathbb{E}}\left[{\bf X}_{t}{\bf X}_{t}^{\prime}\right]=\frac{\tau_{x}^{2}}{1-(\phi_{x}^{T})^{2}}\bm{\Omega}_{x,pcar}^{-1}$
indicates a zero-lag covariance matrix, and $\bm{\Gamma}_{x,ar}$ is the correlation matrix of a first-order autoregressive process with generic element $\left(\bm{\Gamma}_{x,ar}\right)_{jk}=\left({\phi_{x}^{T}}\right)^{|j-k|}$. The spatial dependence is defined by the matrix $\bm{\Omega}_{x,pcar}$, which has a proper conditional autoregressive (PCAR) specification (Banerjee et al., 2014, Besag, 1974, Sahu, 2022):
$\bm{\Omega}_{x,pcar}={\bf D}-\phi_{x}^{S}{\bf W}\,,$ (12)
where ${\bf D}$ is a diagonal matrix with the number of neighbors of each HD on the diagonal, ${\bf W}$ is the binary adjacency matrix, and $\phi_{x}^{S}\in(0,1)$ is the spatial autocorrelation parameter. ${\bf W}$ is an ${N}\times{N}$ matrix with entries $w_{ij}=1$ if areal units $i$ and $j$ share a border, and $w_{ij}=0$ otherwise; also, $w_{ii}=0$.
We set $\rho_{xz}=0.5$, $\tau_{x}^{2}=4\times 10^{-4}$, and $\tau_{z}^{2}=1\times 10^{-3}$, and then use different values for the spatial and temporal autocorrelation parameters to define the scenarios in Table 1 of the main manuscript.
Once all the matrices in Equation (11) are defined, we simulate the exposure from its marginal distribution, then the unmeasured confounder from the conditional distribution ${\bf Z}|{\bf X}\sim\mathcal{N}\left(\bm{\mu}_{z|x},\bm{\varSigma}_{z|x}\right)$ where:
$\bm{\mu}_{z|x}$ $={\rm\mathbb{E}}\left[{\bf Z}|\mathbf{X=x}\right]=\bm{\mu}_{z}+\bm{\varSigma}_{zx}\bm{\varSigma}_{x}^{-1}(\mathbf{x}-\bm{\mu}_{x}),$
$\bm{\varSigma}_{z|x}$ $={\rm\mathbb{V}ar}\left[\mathbf{Z}|\mathbf{X=x}\right]=\bm{\varSigma}_{z}-\bm{\varSigma}_{zx}\bm{\varSigma}_{x}^{-1}\bm{\varSigma}_{xz}\,.$
Finally, the outcome is generated using the following model:
$Y_{it}|\mu_{it}$ $\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{P}oi(\mu_{it})$ (13)
$\vartheta_{it}$ $={\color[rgb]{0,0,0}o_{it}+\beta_{0}+\beta_{1it}X_{it}+Z_{it}+u_{it},\quad u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})}\,,$ (14)
where $\mu_{it}=\exp(\vartheta_{it})$, $o_{it}=\log(E_{it})$ is the offset term, and $\tau_{u}^{2}=10^{-3}$. The expected counts, $E_{it}$, are obtained by using the resident population aged $\geq 65$ years and the age-specific standardized mortality rates of year 2018. The population and the rates are assumed to be constant in time.
E.2 Performance Assessment
To complete the results of Simulation Study 1, Tables E.4E.6 report, for each scenario and model, the value for the three indices, namely:
  • average mean squared error, $MSE_{avg}=\frac{1}{{N}{T}}\sum_{i=1}^{{N}}\sum_{t=1}^{{T}}\left[\frac{1}{100}\sum_{j=1}^{100}\left({\rm\mathbb{E}}^{(j)}[P_{it}|\bullet]-P_{it}\right)^{2}\right]$, where ${\rm\mathbb{E}}^{(j)}[P_{it}|\bullet]$ is the posterior mean for the $j$-th simulated dataset;
  • maximum absolute bias, $MAB=\max_{i,t}\frac{1}{100}\sum_{j=1}^{100}|{\rm\mathbb{E}}^{(j)}[P_{it}|\bullet]-P_{it}|$
  • average mean absolute error, $MAE_{avg}=\frac{1}{{N}{T}}\sum_{i=1}^{{N}}\sum_{t=1}^{{T}}\left[\frac{1}{100}\sum_{i=1}^{100}|{\rm\mathbb{E}}^{(j)}[P_{it}|\bullet]-P_{it}|\right]$
Here, the differences are computed with respect to $P_{it}=100(\exp\{10(\beta_{1it})\}-1)$, i.e. the true simulated percent change of risk, which varies spatio-temporally.
Tables E.4E.6 indicate that, SDGLMC is the best model among those compared in terms of $MSE_{avg}$, $MAB$, and $MAE_{avg}$ for all scenarios. Recalling the decomposition of $\beta_{1it}$ in Section 3.1 of the main manuscript, we have also computed the average mean squared error for each of the components of the exposure effect – global, temporal and spatial. SDGLMC provides the lowest values in most of the cases, and the improvements over the other approaches are more pronounced for the temporal component.
Since we want the overall error to be as low as possible, and since the shape of the exposure effect is not known in practice, simulation results suggest that the proposed model is the most robust choice across all scenarios.
Table E.4: Simulation Study 1: $MSE_{avg}$ calculated for each model considered under each Scenario. The lowest values are indicated in italic.
Scenario Component Formula SDGLMC Null GLMadj Dummy Periodic JDZ
S1 Overall $100(\exp\{10\beta_{1it}\}-1)$ 0.669 36.214 2.417 1.888 2.780 2.522
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.115 34.568 0.193 0.142 0.225 0.275
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.200 1.418 1.418 0.889 1.705 1.418
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.354 0.228 0.806 0.855 0.844 0.829
S2 Overall $100(\exp\{10\beta_{1it}\}-1)$ 5.617 20.440 6.596 6.022 8.118 6.681
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 5.246 13.331 4.807 4.897 6.568 4.893
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.103 1.418 1.418 0.837 1.204 1.418
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.268 5.691 0.371 0.277 0.288 0.369
S3 Overall $100(\exp\{10\beta_{1it}\}-1)$ 0.857 31.247 2.286 1.699 2.222 2.337
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.509 29.444 0.370 0.310 0.423 0.417
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.133 1.418 1.418 0.886 1.281 1.418
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.216 0.385 0.498 0.500 0.506 0.502
Table E.5: Simulation Study 1: $MAB$ calculated for each model considered under each Scenario. The lowest values are indicated in italic.
Scenario Component Formula SDGLMC Null GLMadj Dummy Periodic JDZ
S1 Overall $100(\exp\{10\beta_{1it}\}-1)$ 1.212 7.770 3.816 3.796 4.010 3.618
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.295 5.856 0.189 0.022 0.340 0.316
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.276 2.100 2.100 1.913 2.497 2.100
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.934 1.373 1.859 1.873 1.787 1.788
S2 Overall $100(\exp\{10\beta_{1it}\}-1)$ 2.898 5.896 3.958 4.048 5.117 3.904
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 2.280 3.647 2.183 2.204 2.555 2.203
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.573 2.100 2.100 2.003 2.190 2.100
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 1.198 1.095 0.387 0.356 0.372 0.313
S3 Overall $100(\exp\{10\beta_{1it}\}-1)$ 1.506 7.683 2.767 2.769 3.607 2.791
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.599 5.413 0.479 0.403 0.529 0.523
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.211 2.100 2.100 1.968 2.250 2.100
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.740 1.691 1.100 1.024 1.114 1.168
Table E.6: Simulation Study 1: $MAE_{avg}$ calculated for each model considered under each Scenario. The lowest values are indicated in italic.
Scenario Component Formula SDGLMC Null GLMadj Dummy Periodic JDZ
S1 Overall $100(\exp\{10\beta_{1it}\}-1)$ 0.654 5.856 1.287 1.111 1.330 1.315
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.297 5.856 0.340 0.286 0.400 0.422
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.355 1.065 1.065 0.801 1.046 1.065
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.468 0.374 0.728 0.751 0.744 0.739
S2 Overall $100(\exp\{10\beta_{1it}\}-1)$ 2.281 3.868 2.226 2.213 2.564 2.242
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 2.280 3.647 2.183 2.204 2.555 2.203
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.253 1.065 1.065 0.794 0.823 1.065
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.402 1.908 0.485 0.415 0.425 0.485
S3 Overall $100(\exp\{10\beta_{1it}\}-1)$ 0.735 5.413 1.272 1.060 1.176 1.285
Global $100(\exp\{10\overline{\delta}_{1}\}-1)$ 0.599 5.413 0.479 0.404 0.529 0.523
Temporal $100(\exp\{10\delta_{1t}\}-1)$ 0.287 1.065 1.065 0.806 0.907 1.065
Spatial $100(\exp\{10\delta_{1i}\}-1)$ 0.365 0.495 0.557 0.556 0.561 0.558
Appendix F Simulation Study 2
We now illustrate Simulation Study 2, where we want to assess whether SDGLMC is able to recover the first derivative of the unknown functional form, which represents the marginal effect of the exposure on the outcome. To this end, we consider the Scenario S1 again, but we assume that the linear predictor is as follows:
$\vartheta_{it}=o_{it}+\beta_{0}+f(X_{it};\bm{\gamma}_{f})+Z_{it}+u_{it}\,,$
where $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}+\gamma_{f2}X_{it}^{2}+\gamma_{f3}X_{it}^{3}$, $\gamma_{f0}=-0.08$, $\gamma_{f1}=0.02$, $\gamma_{f2}=-8\times 10^{-5}$, $\gamma_{f3}=5\times 10^{-8}$. Henceforth, the true values of the time-varying coefficients are $\beta_{1t}=0.02-1.6\times 10^{-4}\widehat{X}_{it}+1.5\times 10^{-7}\widehat{X}_{it}^{2}$, and $\beta_{0it}=\beta_{0}-0.08+0.02\widehat{X}_{it}-8\times 10^{-5}\widehat{X}_{it}^{2}+5\times 10^{-8}\widehat{X}_{it}^{3}+Z_{it}$. All the other parameters are the same as in the previous Section.
Table F.7 reports the $MSE_{avg}$, $MAB$, and $MAE_{avg}$ that were obtained. The small values of the indices indicate that the proposed model is able to recover the first derivative of the unknown functional form. Some confounding bias is still present, but it is not attributable to the Taylor’s series approximation, since it is not present when the data are generated without unmeasured confounder (i.e., $Z_{it}=0$).
Table F.7: Simulation Study 2: $MSE_{avg}$, $MAB$, and $MAE_{avg}$ calculated for SDGLMC and SDGLMC_single under Scenario S1 with $f(X_{it};\bm{\gamma}_{f})=\gamma_{f0}+\gamma_{f1}X_{it}+\gamma_{f2}X_{it}^{2}+\gamma_{f3}X_{it}^{3}$. The lowest values for each index are indicated in italic.
$MSE_{avg}$ $MAB$ $MAE_{avg}$
Component SDGLMC SDGLMC_single SDGLMC SDGLMC_single SDGLMC SDGLMC_single
Overall 1.381 1.897 2.288 3.330 0.929 1.094
Global 0.050 0.398 0.170 0.614 0.185 0.614
Temporal 0.376 0.503 2.000 2.579 0.481 0.544
Spatial 0.945 0.984 0.452 0.860 0.762 0.741
F.1 On The Removal of Large-Scale Temporal Trend.
We argue that the term $X_{it}-\widehat{X}_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$, should be used in place of simply $X_{it}$ in the DGLM. Removing the overall temporal trend from the exposure not only follows from theoretical arguments (see Equation (3) in the main paper), but also it is crucial to control for confounding bias. The local variations in the exposure are in fact less likely to be confounded by the unmeasured confounder, which is assumed to be spatially and temporally smooth (Janes et al., 2007, Greven et al., 2011) – see also Section 3 of the manuscript. To illustrate this with simulated data, we compare the proposed SDGLMC with a reparameterized model that does not remove the overall trend from the exposure. Essentially, if the first-order Taylor’s series is evaluated at 00, Equation (6) in the paper can be rewritten as:
${\color[rgb]{0,0,0}\vartheta_{it}=o_{it}+\beta_{0it}^{*}+\beta_{1t}^{*}X_{it}+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+\tilde{u}_{it}\,,}$
where $\beta_{1it}^{*}=f^{\prime}(0;\bm{\gamma}_{f})$ and $\beta_{0it}^{*}=\beta_{0}+f(0;\bm{\gamma}_{f})+Z_{it}$.
We refer to this model as SDGLMC_single, since it is a single-step approach that does not require to estimate the overall spatio-temporal trend.
The results for the reparameterized approach are reported in Table F.7, where it is confirmed that SDGLMC outperforms SDGLMC_single according to all indices.
Appendix G Application: Further Analyses
G.1 Comparison with Restricted Model
We assess the impact that imposing some constraints on the proposed approach has on the estimated exposure effect. Specifically, we remove the measured confounders from the model specification. This is equivalent to imposing that $\bm{\alpha}={\bf 0}$ in Equation (7) of the main text. Hence, the linear predictor is specified as follows:
${\color[rgb]{0,0,0}\bm{\vartheta}_{t}=\mathbf{o}_{t}+{\bf 1}_{N}(\overline{\delta}_{0}+\delta_{0t})+\tilde{\bm{\delta}}_{0}+\bm{\delta}_{0t}^{*}+\mathbf{\tilde{X}}_{t}(\overline{\delta}_{1}+\delta_{1t})+\mathbf{\tilde{X}}_{t}\odot\tilde{\bm{\delta}}_{1}+\tilde{{\bf u}}_{t}\,.}$
We refer to the implied model as SDGLMCX to emphasize that the exposure is the only covariate. Figure G.1 shows the posterior mean of the percent change of risk, as defined above, under the SDGLMCX (in green) and the proposed DGLM (in red). It suggests that, in general, the temporal dynamic of the exposure effect would not be affected much if the measured confounders were not included in the model. Henceforth, the proposed approach can provide accurate estimates even in complex scenarios.
However, it is worth noting that SDGLMCX provides greater estimates than SDGLMC during the summer, while during the winter the estimates are quite similar. This suggests that the confounding effect may vary across time. We hypothesize that this is due to air temperature exceeding a certain threshold; in other words, confounding may occur due to heat waves. In fact, heat waves are demonstrated to have an increased effect on mortality during summer days with relatively high levels of PM2.5 concentrations, and this positive correlation may be the source of the overestimation (Analitis et al., 2008, 2014). Another study observed a similar result for the population aged 75 years or more (Analitis et al., 2018). Another possible explanation for the overestimation could be the modification of the PM2.5 chemical composition, with high levels of the most toxic compounds developing in summer (Vinson et al., 2020).
Figure G.1: Estimated time-varying percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations. The red line represents the posterior mean of SDGLMC, while the green line the posterior mean of SDGLMCX.
G.2 Comparison with a GLM with interactions
We argue that the summer peaks of the exposure effect are likely due to potential interactions between PM2.5 and temperature or other confounders. To provide further evidence, we compare the proposed approach with a model that includes an interaction term between PM2.5and each measured confounder. Therefore, we consider a hierarchical model, referred to as GLMint, that follows the same structure as defined in Section D, but with $\mathbf{A}_{it}=(X_{it},X_{it}{{\bf M}}_{it}^{\prime})^{\prime}$ and $\mathbf{C}_{it}=(1,{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$.
Given the hierarchical structure of GLMint, the exposure effect can be estimated for each HD by summing the PM2.5 coefficient and the interaction coefficients (multiplied by the confounder values measured in the respective HD at each point in time). The estimates are then averaged across the HDs to obtain a single estimate for each time point.
Figure G.2 shows the posterior mean of the percent change of risk (as defined in previous sections) under SDGLMC (in red) and GLMint (in dark blue). The estimates from both models appear similar during winter. However, GLMint may overestimate the exposure effect during summer, with peaks approximately equal to $16\%$, which is unexpected for short-term associations. This suggests that GLMint is not able to capture the temporal dynamics of the exposure effect as well as SDGLMC.
Given that the effect estimated by GLMint is not smooth in time, another hierarchical model, GLMint_smooth, is added to the comparison. It includes interaction terms between PM2.5 and smoothed versions of temperature and relative humidity, and is depicted in light blue in Figure G.2. The smoothed confounders are obtained by using a principal spline with $15$ spatial bases and $15$ temporal bases. The results show that the temporal dynamics under GLMint_smooth and SDGLMC are similar in 2018 and 2022. However, it seems that GLMint_smooth could underestimate the effect in 2019, and could overestimate it during the pandemic. Furthemore, GLMint_smooth and the Periodic model seem to provide similar estimates (see Figure 3 in the main manuscript).
Finally, this analysis corroborates the hypothesis of effect modification by unmeasured spatio-temporal confounders, which is not well captured by GLMint or GLMint_smooth. The results indicate that the possibility of a more complex interaction between PM2.5 and the confounders cannot be ruled out.
Figure G.2: Estimated time-varying percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations.